OpenStudy (anonymous):

Log subscript 4(4x) +log subscript 4(x)= log subscript 4(64) solve for x

OpenStudy (loser66):

\[\log_{4} 4x+\log_{4}x=\log_{4} 64?\]

OpenStudy (anonymous):

yes

OpenStudy (loser66):

log a +log b = log (a*b) apply that from LHS, tell me what you have?

OpenStudy (loser66):

@shaq2341 I guide you only, you have to work. I don't work for you

OpenStudy (anonymous):

okay so i got x=4

OpenStudy (loser66):

how to get this?

OpenStudy (anonymous):

since the bases were the same , the formula came down to 4x*x=64

OpenStudy (anonymous):

4x^2=64 . Divide by 4 , then x^2=16 which is equal to 4.

OpenStudy (loser66):

thanks for teaching me back, I got the answer from yours , bingo, give you medal for being my tutor. hehehe...